what is the area of a rectangle length 8 meters perimeter 30 meters

Answer by checkley77(12844)

L=8 METERS IS THE LENGTH.
2L+2W=30
2*8+2W=30
16+2W=30
2W=30-16
2W=14
W=14/2
W=7 METERS IS THE WIDTH.
AREA=6*8=48 ANS. FOR THE AREA.